Restart Night #7 28 Aug 2020 UT

OSA: S. Allanson

ISp: C.Boylard

Mtn Manager: L. Durham

ISA: O. Kuhn (Tucson)

Special Guest Stars: J. Hill (Tucson), B. Rothberg (Tucson)


Summary:

Clouds delayed opening and smoke closed us briefly but, overall, it was a productive night.

The AGW1 alignment went well - the guiding hotspot was not changed at all, though the position of the rotator center on the guide camera had changed because of the encoder replacement. The initial position of SX M3 gave the best alignment/minimum wobble. After John had completed this, we switched to LBCR to follow up on two non-sidereal targets. Passing clouds and moon hampered these observations, but there were a number of observations with stars visible and hopefully the objects were both detected. We switched to binocular MODS to observe the transient and a standard star, and the clouds cleared out by the time the acquisition was done. Next, I confirmed that the blind offset acquisitions with MODS worked well, and watched the PCSGUI as the detxy rel and radec abs offsets were sent and updatepointing was issued.

We ended the night with sets of MODS1 and MODS2 red-only twilight flats at r_sdss and z_sdss.

Issues:

Again, there was significant coma on LBCR (z8). WRS did not correct this but instead sent a large astigmatism correction. When we first collimated with MODS2, there was also large coma (z7 up to 3500 nm).

RB_Science keeps relaunching the ds9 window rather than displaying the image to the existing window.

M1B was out of service due to the blue grating error. There is plan to replace the microLynx controller.

The LBCB cryostat had been removed in the morning for work in the lab.

Not all of the mods data has gone into the archive - we noticed that the mods2 data from 20200822 was not in /Repository/20200822. A note will be added to IT 8160.

Proposed Plan:

  • IF AGW-1 IS RELEASED
    • AGW-1 alignment - (1 hr) - (CRITICAL)
      • DONE correct pointing
      • DONE locate rough guiding hotspot
      • DONE verify SH spots are illuminated
      • DONE verify collimation
      • DONE adjust probe to put rotator center at hotpsot
      • DONE refine guiding hotspot center and update as needed
    • SX M3 Alignment (CRITICAL) - (2 hr)
      • DONE Pupil wobble - requires LFBG/AGW1 to guide, collimate, SH spots well illuminated, and hotspot centered - can be done before AGW Rot ZP and Transform (2 hrs) (high)
    • PEPSI FPU alignment (1 hr) - medium
  • LUCI-2
    • Test Imaging
      • Dither in the field (J1601 until about 10:15pm local time)
      • after 10:15pm TBD
  • MODS-2
    • DONE Transient redshift observation: object available from 04:30UT to sunrise. moon sets 08:00 UT
      • | ZTF20abwysqy | AT2020scz | 6.7855741 | +34.0273043 | g | 20.86 | 0.04 | 2459087.9181366 |
        • 2x900 sec exposures per MODS channel - total time w/overheads 45min
  • LBC-Red
    • Additional NS imaging for space rocks
      • DONE 2017 UR4 --- available from sunset to 07:30 UT
      • DONE 2018 UQ1 --- available from 05:30 UT to sunrise

Log

UT 01:30 I ran up MODS1+2, LBCR and LUCI2 (will run up geirs on LUCI1 next). I tried again to power cycled the microLynx controller 9 for the MODS1 blue grating, but to no avail. I took sieveSnaps with both MODS and am taking some biases to check LBCR. There is a problem with RB_Science in that it keeps relaunching the ds9 window, rather than simply loading the new images into the window. I don't think anything has changed and this may be a timing issue - just a guess. It's annoying because the ds9 will come up on the current workspace and does not stay on the one on which it was launched.

Open

UT 02:45 Steve is opening. The clouds that were overhead at sunset are dissipating.

AGW1 alignment

UT 02:55 Pointing done.

UT 02:58 Finding the hotspot - ACTIVE preset to a pointing star - The old hotspot already has us in the pinhole!

UT 03:04 Collimation - ACTIVE preset to r=12 (probe is far from rotator center as PA changes pointing) - pupil position is OK, seeing is 1.2"

following BGTechProcedures

image files at /lbt/data/share/tcs/lbto.20200828/ as seen from obs4

UT 03:12 gcsclient readGuideCam left -e 80000 ---> guide 0001
208,92 = 416,184 = 431,223 (want 249,196 = old hotspot)
decrease X by 167 pix = 8.35 arcsec = 5.0 mm = increase ROT.homeoffset by 340 counts (from 24557 to 24898)

UT 03:26 back to pointing star with ACQUIRE preset IE=-4 CA=40

UT 03:28 ACTIVE preset to r=13.5 AO681 - seeing is 0.7

UT 03:30 ACTIVE preset to r=16 - The old hotspot is confirmed to be good which is remarkable given all the work we did on AGw1.

UT 03:36 gcsclient readGuideCam left -e 80000 ---> guide 0002
141.5,84.1 = 283,168.1 = 298,207.1
decrease X by 49 pixels = 2.45 arcsec = 1.47 mm = increase homeoffset by 100 counts
decrease Y by 11 pixels = 0.55 arcsec = 330 microns = increase ooffset by 300 microns
new values homeoffset = 24998 ooffset = 419562

UT 03:52 Close for smoke

UT 04:29 Re-opening as smoke smell is much fainter and there seems to be no recent accumulation and the flashlight beam is clear.

SX M3 alignment by measuring Pupil Wobble

procedure in ComTools

UT 04:39 ACTIVE preset to r=13.5 AO681 - seeing is 1.1 arcsec

UT 04:47 stop wfs

image files at /lbt/data/share/tcs/lbto.20200828/ as seen from obs4

analysis files in /home/lbto/data/20200828_lfbg/
Where did all the old data go? It is still there - John got confused between /home/lbto/data/ and /home/lbto/Desktop/data/ for a while.

Starting M3 position: TIPabs = -180, TILTabs = -210

UT 04:48 pupil_collect
cat pupil_04:47:12.wobble
# Shack-Hartmann Pupil Wobble data
# x_center   y_center   rotangle    imagename
#----------------------------------------------
    378.71    242.06    410.86   left_wfsc000001_dd.fits
    377.86    245.87    351.53   left_wfsc000002_dd.fits
    379.60    247.53    292.18   left_wfsc000003_dd.fits
    382.74    246.38    233.49   left_wfsc000004_dd.fits

+50 Tip 0 Tilt dIE +1.5 dCA +4
We spent about 5 minutes understanding why the required corrections didn't match the wiki page. Steve corrected the wiki page by adding the missing brackets to the equations. It would have been faster to use supportscripts/MetrologySupport/m3tt.py which automatically applies the IE/CA compensation.

UT 05:11 pupil_collect
cat pupil_05:11:04.wobble
# Shack-Hartmann Pupil Wobble data
# x_center   y_center   rotangle    imagename
#----------------------------------------------
    379.80    238.26    422.44   left_wfsc000005_dd.fits
    376.40    244.75    363.02   left_wfsc000006_dd.fits
    378.79    248.03    303.61   left_wfsc000007_dd.fits
    382.91    249.00    244.63   left_wfsc000008_dd.fits

+50 Tip +50 Tilt dIE=-2.7 dCA=+1.2

UT 05:xx pupil_collect
 cat pupil_05:27:42.wobble
# Shack-Hartmann Pupil Wobble data
# x_center   y_center   rotangle    imagename
#----------------------------------------------
    379.52    239.73    428.99   left_wfsc000009_dd.fits
    374.63    242.85    369.45   left_wfsc000010_dd.fits
    378.00    248.38    309.94   left_wfsc000011_dd.fits
    382.86    248.87    250.77   left_wfsc000012_dd.fits

First Tip made it better (wrong). First Tilt made it worse. Put Tilt back to 0, Double Tip to +50.

+100 Tip 0 Tilt

UT 05:42 pupil_collect (some clouds going by)
cat pupil_05:41:27.wobble
# Shack-Hartmann Pupil Wobble data
# x_center   y_center   rotangle    imagename
#----------------------------------------------
    380.12    238.51    433.70   left_wfsc000013_dd.fits
    374.27    240.97    374.11   left_wfsc000014_dd.fits
    377.58    248.32    314.58   left_wfsc000015_dd.fits
    381.70    250.25    255.32   left_wfsc000016_dd.fits

-50 Tip 0 Tilt - dIE=-5 dCA=-12

UT 05:58 pupil_collect
cat pupil_05:58:07.wobble
# Shack-Hartmann Pupil Wobble data
# x_center   y_center   rotangle    imagename
#----------------------------------------------
    380.70    240.77    438.74   left_wfsc000017_dd.fits
    376.69    242.37    379.09   left_wfsc000018_dd.fits
    378.80    247.41    319.44   left_wfsc000019_dd.fits
    382.14    248.54    260.11   left_wfsc000020_dd.fits

Summary Statistics
pupil_05:58:07.wobble 379.5825 244.7725 7.72728606951 4
pupil_05:41:27.wobble 378.4175 244.5125 11.2947664872 4
pupil_05:27:42.wobble 378.7525 244.9575 9.68562594776 4
pupil_05:11:04.wobble 379.475 245.01 9.61770762708 4
pupil_04:47:12.wobble 379.7275 245.46 5.52029664058 4

  • SX M3 Pupil Wobble Plot: (world coords adjusted in pupil_wobble.par)
    Screenshot at 2020-09-01 03-17-17.png

Conclusion is that the initial tip-tilt settings (0,0 relative) are best. Maybe there is a slight leaning toward -10 Tip?

LBCR

UT 6:13 switching to MODS1/LBCR

UT 06:34 Slewing to copointing star for first NS target.
  • dohybrid, /redonly
    • There's a lot of coma initially, see lbcr.20200828.063934.fits. dohybrid is not removing it (it is instead correcting astigmatism, z6). The first FPIA corrections again include a large Z8.
    • We should edit the configuration file for WRS so that WRS does not send astigmatism, only sends coma.
  • ptautoadjust for pointing correction
    • The CA and IE offsets were 16.3" -12.5"
  • Observing the field 2017 UR4
    • 065852-070540 9pt dither with 60-sec exposures at each dither position.
    • object is pretty faint (23.63) so we are taking a series of 3 150-sec exposures at each of 2 dither positions. Clouds are also coming over where we are looking.
      • 071028 - saturating. Stopped the script and created one with a set of 3 60-sec exposures. The clouds & moon make the background too high for 150-sec. The guiding corrections were looking a bit odd for at least one - but maybe the conditions played a role?
      • 3x60-sec: starting at 072227
  • dofpia, /redonly on field of 2018 UQ1
  • ptautoadjust on CP star, 074259
    • CA and IE offsets are -13" and 7"
  • Observing the field of 2018 UQ1
    • first script took 2 60-sec exposures at 0,0 and 30",30": 075121, 075303
    • second script is taking 3 180-sec exposures, but the cloud are coming and some of these look like skyflats.
      • first position (0,0): 080419,080758, 081136
      • second position (-4", -8"): 081514, 081852, 082232

Bino-MODS

UT 08:27 - Steve is reconfiguring for binocular MODS. On the right side, 3500nm of Z7.

UT 08:49 GRB observation
  • acqBinoMODS AT2020scz_Red_Acq_M1.txt AT2020scz_Blue_Acq_M2.txt
    • mods1r: acq images are 2 and 3
      • modsAlign -y -8
      • -0.747 -7.828 rel
    • mods2b: acq images are 2 and 3
      • modsAlign -y -8
      • 4.002 -10.038 rel
    • confirmatory thru-slit images:
      • mods1r and mods2b 4: both look pretty good - we see the object through the slit
  • execBinoMODS AT2020scz_Red_Obs_M1.txt AT2020scz_Blue_Obs_M2.txt
    • During the first set of exposures: m1r 5,6 and m2b 5,6, m2r 2,3; the object is transiting at 88:30 elevation, but guiding is holding on. The guide star is lagging a bit and is off the hotspot by about 0.5" at times. The seeing is about 0.6"-0.7". There are some emission lines, it seems.
    • Taking a 2 more exposures. mods1r 7,8 and mods2b 7,8, mods2r 4,5
UT 10:22 Feige110
  • acqBinoMODS feige110.acq
    • MODS2 timed out (120-sec) in changing AGWfilter. Must have been a comm glitch. Retry worked.
    • a pointing correction would have been needed, but as the object is at low airmass and setting, we decided to go to another star.
UT 10:34 G191B2B
  • acqBinoMODS g191b2b.acq
    • a pointing correction was needed
    • imcslock failed on mods2, a retry worked
    • red channel acq images will be m1r 9 and m2r 6
    • mods1 offset: 0.453 11.368
    • mods2 offset: 4.783 7.634
    • confirmatory thru-slit images are m1r 10 and m2r 7. Star is on the right side for both mods1 and mods2. Adjusting...
      • mods1r: -0.638 -0.043. Next image, 11, looks well-centered.
      • mods2r - image 0007 has the mask IN, but the header reads STOW, so modsAlign is not auto-calculating the offset needed. Could edit header - but taking another image, instead. Based on mods2r 0008, an offset of -0.947", -0.091" is needed to center the object in the slit. Next image, 0009, looks good.
  • execBinoMODS g191b2b_M1RedGrating.obs g191b2b_M2DualGrating.obs

Checking blind offset acquisitions on Stone A

  • acqBinoMODS StoneA _9to13.txt
    • modsAlign -y 11 mods1r.20200828.0015.fits 16 => offsetxy -0.869 11.326 rel
    • modsAlign -y 9 mods2r.20200828.0013.fits mods2r.20200828.0014fits => offsetxy 3.515 7.508 rel
    • I watched PCSGUI as offset DETXY was populated with the offsets and then zero'd out as the updatepointing command was issued.
    • confirmatory thru-slit images mods1r 17 and mods2r 15, look very good, but to insure that the tweaks are handled properly, I send +5 detxy y to left and -1" detxy x to right side. I do not absorb and then hit return to send the blind offset. Immediately the detxy offsets are zero'ed out, as expected, and then the radec offset is sent. A screenshot of the PCS GUI is attached which shows the offset RADEC holding the blind offsets and the achieved positions and target base differing by what looks like the offset ra.
    • Images m1r 19 and m2r 17 are of the target. By eye, the positions of the target appear to coincide nicely with those of the reference star.
  • acqBinoMODS StoneA _23to17.txt
    • uses the 1" slit
    • modsAlign -y 11 ... offsets are -0.627 11.795 rel. Once absorbed, these are added to the Pointing Origin Base.
    • modsAlign -y 9 ... offsets are 3.665 7.880 rel
    • Images m1r 23 and m2r 21 are of target. It looks well centered.

Twilight sky flats

  • Red-only
  • Blank 03+33
  • r _sdss 10-sec m1r: 26-29 and m2r: 24-28. These have count levels of 2000,3000 up to 12,000. For MODS1, there are some problems cascading from having cntl'cd a sync preset but not asking Steve to cancel the preset. I eventually aborted the MODS1 script as it got into timing errors. A 2nd set: m1r:29-x has 17k - xxk and m2r:29-xx has 27k - 32k
  • z_sdss mods1 set needs to be picked through but the mods2 set is pretty good. Both are image numbers 34-38
  • We got good sets of red-only r_sdss and z_sdss twilight flats on both MODS.
UT 12:38 Steve is closing up.

UT 12:59 I am taking a set of 25 LBCR biases and 10 MODS 3K and 1K biases.

UT 13:40 I put both MODS to sleep and turned off LBCR and disconnected from the TCS.

-- BarryRothberg - 27 Aug 2020

Topic attachments
I AttachmentSorted ascending Action Size Date Who Comment
Screen Shot 2020-08-28 at 4.35.36 AM.pngpng Screen Shot 2020-08-28 at 4.35.36 AM.png manage 144 K 28 Aug 2020 - 11:42 OlgaKuhn Snapshot of PCSGUI after the blind offset was made
Screenshot at 2020-09-01 03-17-17.pngpng Screenshot at 2020-09-01 03-17-17.png manage 46 K 01 Sep 2020 - 03:23 JohnHill SX M3 Pupil Wobble Plot
Topic revision: r18 - 01 Sep 2020, JohnHill
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback